OpenStudy (anonymous):

The square root of 64 plus the square root of 121 equal?

OpenStudy (anonymous):

19

OpenStudy (anonymous):

8+11

OpenStudy (anonymous):

19? Sqrt of 121= 11, Sqrt of 64= 8. 11+8

OpenStudy (anonymous):

19

OpenStudy (anonymous):

=\[\sqrt{64} +\sqrt{121}\] =8+11 =19 PLEASE CLICK THE GOOD MEDAL BUTTON TNX